OpenStudy (rainbow_dashie):

Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct.

OpenStudy (rainbow_dashie):

A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. is a positive integer

OpenStudy (bibby):

is D p/q?

OpenStudy (rainbow_dashie):

d. is a positive integer

OpenStudy (bibby):

that doesn't make sense yo

OpenStudy (bibby):

let's assign p and q numbers p=3 and q=5

OpenStudy (pottersheep):

Bibby is right, your choice for D doesn't make sense

OpenStudy (bibby):

is 3+5 an integer? 3-5? 3*5? 3\(\div\)5?

OpenStudy (pottersheep):

A. p + q is a positive integer TRUE! Two positive numbers added = a positive number! Always! B. p – q is a negative integer Maybe not. Ie if P = 10 and q = 1, p-q = 10-1 = +9. FALSE C. p • q is a positive integer Yes. Two positive numbers multiplied together is still a positive integer D. is a positive integer (incomplete choice)

OpenStudy (rainbow_dashie):

IK

OpenStudy (bibby):

good job, I slipped up on my number choice for b logically d is division btw @pottersheep

OpenStudy (rainbow_dashie):

thx much!!!

OpenStudy (rainbow_dashie):

Is the set of multiples of 5 closed under division? Explain why or provide a counterexample if not.

OpenStudy (rainbow_dashie):

A. Yes, because the quotient of any two multiples of 5 is also a multiple of 5. B. Yes, because the quotient of any two multiples of 5 is also a multiple of 10. C. No, and a counterexample is 100 ÷ 5 = 20. D. No, and a counterexample is 25 ÷ 50 = .

OpenStudy (rainbow_dashie):

c and d. /